# spacing=nonfrench

### v1

@@ Abstract
Large Language Models (LLMs) have advanced code generation and software automation, but are fundamentally constrained by limited inference-time context and lack of explicit code structure reasoning. We introduce Kodezi Chronos [51], a next-generation architecture for autonomous code understanding, debugging, and maintenance, designed to operate across ultra-long contexts comprising entire codebases, histories, and documentation—all without fixed window limits.
Kodezi Chronos leverages a multi-level embedding memory engine, combining vector and graph-based indexing with continuous code-aware retrieval. This enables efficient and accurate reasoning over millions of lines of code, supporting repository-scale comprehension, multi-file refactoring, and real-time self-healing actions.
Our evaluation introduces a novel Multi Random Retrieval benchmark, specifically tailored to the software engineering domain. Unlike classical retrieval benchmarks, this method requires the model to resolve arbitrarily distant and obfuscated associations across code artifacts, simulating realistic tasks such as variable tracing, dependency migration, and semantic bug localization. Chronos outperforms prior LLMs and code models—demonstrating a 23% improvement in real-world bug detection and reducing debugging cycles by up to 40% compared to traditional sequence-based approaches.
By natively interfacing with IDEs and CI/CD workflows, Chronos enables seamless, autonomous software maintenance, elevating code reliability and productivity while reducing manual effort. These results mark a critical advance toward self-sustaining, continuously optimized software ecosystems.
@@ I INTRODUCTION
Recent advancements in large language models (LLMs) have transformed code generation, review, and reasoning tasks [1, 2, 10]. However, debugging—the most time-consuming and critical aspect of software development—remains largely unsolved. While tools like GitHub Copilot [11], Cursor, and Claude excel at code completion [12], they fundamentally misunderstand debugging as a multi-faceted, context-heavy process that spans entire repositories, historical commits, CI/CD logs, and runtime behaviors. Production debugging requires reasoning across files separated by thousands of lines [13], understanding temporal code evolution, and correlating seemingly unrelated symptoms to root causes buried deep in dependency chains [14].
@@
Current code assistants fail at debugging for three critical reasons: (1) they are trained primarily on code completion tasks, not debugging workflows [15]; (2) they lack persistent memory of past bugs, fixes, and codebase-specific patterns [16]; and (3) their context windows, even when extended to 100K+ tokens, cannot capture the full debugging context needed for complex, multi-file issues. Recent studies show that even state-of-the-art models like GPT-4, Claude-3, and Gemini-1.5 achieve less than 10% success rates on real-world debugging benchmarks [17, 18], often proposing superficial fixes that fail validation or introduce new regressions [19].
@@
Kodezi Chronos represents a paradigm shift: the first debugging language model developed by Kodezi [51], specifically designed, trained, and optimized for autonomous bug detection, root cause analysis, and validated fix generation. Unlike code completion models that generate syntactically correct but often semantically flawed suggestions, Chronos operates through a continuous debugging loop—proposing fixes, running tests, analyzing failures, and iteratively refining solutions until validation succeeds. Built on a novel architecture combining persistent debug memory, multi-source retrieval (code, logs, traces, PRs), and execution sandboxing, Chronos achieves what no existing model can: true autonomous debugging at repository scale.
@@
Chronos is designed for seamless integration with modern development stacks: it operates as an embedded “AI CTO” within CI/CD pipelines, IDEs, and project management tools. Its proactive, event-driven workflow ensures that debugging, documentation generation, refactoring, and even preventative maintenance actions occur autonomously, guided by deep repository memory rather than manual prompts or brittle heuristics.
@@
This paper presents the architecture, memory system, retrieval mechanism, evaluation methodology, and experimental results that establish Kodezi Chronos as the new frontier for autonomous, ultra-contextual codebase management.
II RELATED WORK AND LIMITATIONS OF TRADITIONAL APPROACHES
@@
The emergence of large-scale neural models for source code processing—such as CodeBERT [46], GraphCodeBERT [4], and CodeT5 [5]—has significantly advanced automatic code synthesis, translation, and review. Many of these models, as well as massive LLMs like GPT-3 and GPT-4 [1, 47], are pre-trained on billions of lines of code paired with natural language, learning rich semantic representations for many programming languages.
Despite impressive gains on benchmark tasks, these approaches are fundamentally bottlenecked by attention-based architectures and fixed-size input windows, typically constraining context to tens of thousands of tokens. Practical developer workflows, by contrast, require tools to reason across entire repositories: files, modules, historic commits, documentation, and evolving build systems, far outstripping standard context limits.
State-of-the-art retrieval-augmented generation (RAG) methods extend practical context by embedding and recalling external documents [8], but remain mostly limited to chunked passage retrieval and lack true compositional, dependency-aware memory. Window expansion techniques—such as models with 100k tokens or more [7]—incur prohibitive compute/memory costs and suffer from diluted attention, leading to information loss and degraded performance as codebase size increases.
Existing benchmarks for long-context reasoning, like the ”Needle in a Haystack” pattern, commonly involve embedding a salient but easily differentiable clue in a sea of noise or repetition. Such evaluations unintentionally rely on the model’s ability to match explicit tokens or unusual patterns, providing only a partial test of true retrieval and reasoning ability. In real-world maintenance, however, code elements are frequently unmarked, distributed, and semantically interdependent, requiring the AI not just to ”find,” but to infer context, resolve relationships, and reason about impact [45].
@@
III ARCHITECTURE OF KODEZI CHRONOS
III-A The Output-Heavy Nature of Debugging
@@ III-A1 Input vs Output Token Distribution
  • Total input: Often ¡10K tokens for most real-world debugging tasks
@@ III-A2 Why Output Quality Trumps Input Size
TABLE I: Input vs output characteristics in debugging tasks.
@@ III-A3 Chronos’s Output-Optimized Architecture
    Iterative Refinement Loop: Rather than single-shot generation, Chronos validates outputs through execution, using test results to refine patches—ensuring output quality over quantity.
@@
III-B Core Architecture Overview
Kodezi Chronos is designed as an autonomous memory-driven intelligence layer for code, operating at scales that span entire enterprise repositories, team histories, and auxiliary knowledge sources. Its architecture consists of three core modules: (i) a persistent Memory Engine for continuous graph-based context construction, (ii) an advanced Retriever that dynamically assembles deep semantic context from code and documentation, and (iii) a transformer-based Code Reasoning Model for synthesis, debugging, and orchestration of software changes.

### v2

@@ Abstract
[…]
On 5,000 real-world scenarios, Chronos achieves 67.3% ± 2.1% fix accuracy versus 14.2% ± 1.3% (Claude 4) and 13.8% ± 1.2% (GPT-4.1), with Cohen’s d=3.87 effect size. The system reduces debugging time by 40% and iterations by 65%. Chronos resolves complex multi-file bugs requiring cross-repository understanding and temporal analysis.
Key limitations include 23.4% success on hardware-dependent bugs and 41.2% on dynamic language issues. Theoretical analysis proves O(k log d) retrieval complexity with convergence guarantees. Human evaluation (N=50) shows 89% preference over baselines. Available Q4 2025 (Kodezi OS) and Q1 2026 (API).
@@ I INTRODUCTION
Recent advancements in large language models (LLMs) have transformed code generation, review, and reasoning tasks [1, 2, 3]. In 2025, frontier models like Claude Opus 4 [4] achieve 72.5% on SWE-bench, GPT-4.1 [5] reaches 54.6%, and DeepSeek V3 [6] demonstrates remarkable efficiency with only \$5.6M training cost. However, debugging, the most time-consuming and critical aspect of software development, remains largely unsolved. While tools like GitHub Copilot [7, 8], Cursor, Windsurf [9], and Claude excel at code completion [10], they fundamentally misunderstand debugging as a multifaceted, context-heavy process that spans entire repositories, historical commits, CI/CD logs, and runtime behaviors. Production debugging requires reasoning across files separated by thousands of lines [11], understanding temporal code evolution, and correlating seemingly unrelated symptoms to root causes buried deep in dependency chains [12].
Why does debugging remain fundamentally unsolved for LLMs? Beyond the surface-level challenges, debugging exposes four core limitations of current architectures. First, hallucination under uncertainty: when LLMs encounter ambiguous error states, they confidently generate plausible-sounding but incorrect fixes, often introducing subtle bugs that pass initial tests but fail in production. Second, absence of persistent memory: each debugging session starts tabula rasa, unable to learn from previous encounters with similar bugs or build institutional knowledge about codebase-specific patterns. Third, rigid token limits: even 1M-token contexts cannot capture the full transitive closure of dependencies in modern software, where a bug’s root cause may span dozens of files connected through deep call chains. Fourth, lack of causal reasoning: LLMs excel at pattern matching but struggle with counterfactual reasoning (”what if this variable were null?”) and temporal causality (”which commit introduced this regression?”), both essential for debugging.
@@
Current code assistants fail at debugging for three critical reasons: (1) they are trained primarily on code completion tasks, not debugging workflows [13]; (2) they lack persistent memory of past bugs, fixes, and codebase-specific patterns [14]; and (3) their context windows, even when extended to 1M tokens (Gemini 2.0) or leveraging advanced RAG techniques like HyDE [15] and FLARE [16], cannot capture the full debugging context needed for complex, multi-file issues. Recent studies show that even state-of-the-art models like GPT-4.1, Claude 4 Opus, and Gemini 2.0 Pro achieve less than 15% success rates on real-world debugging benchmarks (COAST, MLDebugging, MdEval) [17, 18], often proposing superficial fixes that fail validation or introduce new regressions [19].
@@
Kodezi Chronos represents a paradigm shift: the first debugging language model developed by Kodezi [20], specifically designed, trained, and optimized for autonomous bug detection, root cause analysis, and validated fix generation. For more information about the model and benchmarks, visit https://chronos.so/ and https://github.com/kodezi/chronos. Kodezi OS information is available at https://kodezi.com/os. Unlike code completion models that generate syntactically correct but often semantically flawed suggestions, Chronos operates through a continuous debugging loop, proposing fixes, running tests, analyzing failures, and iteratively refining solutions until validation succeeds. Built on a novel architecture combining persistent debug memory, multi-source retrieval (code, logs, traces, PRs), and execution sandboxing, Chronos demonstrates debugging performance that significantly exceeds current models, particularly in repository-scale, multi-file scenarios.
@@
Chronos is designed for seamless integration with modern development stacks: it operates as an embedded autonomous maintenance system within CI/CD pipelines, IDEs, and project management tools. Its proactive, event-driven workflow ensures that debugging, documentation generation, refactoring, and even preventative maintenance actions occur autonomously, guided by deep repository memory rather than manual prompts or brittle heuristics.
@@
This paper presents the architecture, memory system, retrieval mechanism, evaluation methodology, and experimental results that establish Kodezi Chronos as the new frontier for autonomous debugging with full repository context.
Why Chronos Dominates 87.1% debugging success vs. 22.9% for GPT-4.1 + SWE-agent 3.8x faster bug resolution through AGR’s precision retrieval 67% fewer false positives via Persistent Debug Memory (PDM) First to handle cross-file, temporal, and runtime-dependent bugs
Our contributions are as follows:
  1. 1.
    We introduce Chronos, a debugging-specific language model architecture integrating persistent memory, adaptive graph-based retrieval (AGR), and a multi-stage fix orchestration loop.
  2. 2.
    We design AGR, a multi-hop, edge-weighted graph traversal system for deep context retrieval tailored to codebases, achieving 92% precision at 85% recall on debugging queries.
  3. 3.
    We construct new debugging-specific evaluation benchmarks, including cycle-aware test loops, retrieval accuracy, and root-cause localization across 12,500 real-world bugs.
  4. 4.
    We demonstrate significant performance gains over Claude 4 Opus, GPT-4.1, and specialized tools, achieving 65.3% debugging success rate on our comprehensive benchmarks (4-5x improvement over GPT-4.1, Claude 4 Opus).
  5. 5.
    We provide comprehensive ablation studies showing each component’s contribution, with detailed case studies and quantitative analysis demonstrating 4-5x improvement over state-of-the-art models.
II RELATED WORK: LIMITATIONS OF EXISTING APPROACHES IN DEBUGGING
II-A Retrieval-Augmented Code Generation
@@
The emergence of large-scale neural models for source code processing, such as CodeBERT [21], GraphCodeBERT [22], and CodeT5 [23], has significantly advanced automatic code synthesis, translation, and review. CodeT [24] introduced code generation with testing, improving reliability through execution feedback. Many of these models, as well as massive LLMs from the GPT-4 family [1, 5], are pre-trained on billions of lines of code paired with natural language, learning rich semantic representations for many programming languages.
State-of-the-art retrieval-augmented generation (RAG) methods have evolved significantly in 2025. Advanced techniques like HyDE (Hypothetical Document Embeddings) [15] generate synthetic documents to improve retrieval quality, Self-RAG [25] uses reflection tokens for dynamic retrieval decisions, and FLARE [16] implements forward-looking active retrieval based on generation confidence. GraphRAG [26] integrates knowledge graphs for structured retrieval, representing a step toward graph-based understanding but still operating with static graphs that lack the dynamic updates required for debugging workflows. Recent work on code-specific retrieval [27] shows promise but lacks the multi-hop reasoning capabilities essential for debugging.
LangChain [28] and DSPy [29] provide frameworks for building complex LLM applications with retrieval components. While these frameworks enable sophisticated pipelines, they fundamentally rely on stateless retrieval without the persistent memory or specialized debugging knowledge that Chronos provides. DSPy’s optimization of prompting strategies and LangChain’s chain-of-thought orchestration, while powerful for general tasks, lack the domain-specific understanding of code dependencies and bug patterns that debugging requires.
Despite impressive gains on benchmark tasks, these approaches are fundamentally bottlenecked by attention-based architectures and fixed-size input windows, typically constraining context to tens of thousands of tokens. Window expansion techniques, such as models with 1M+ tokens (Gemini 2.5, GPT-4.1) [30, 5], incur prohibitive compute/memory costs and suffer from diluted attention, leading to information loss and degraded performance as codebase size increases.
II-B Program Repair and Bug Localization
Recent systems targeting debugging include SWE-agent [13], which achieves 12.3% on the SWE-bench dataset through structured agent-computer interfaces, and AutoCodeRover [17], reaching 22.9% via program structure-aware retrieval. These systems represent significant progress but still operate without persistent memory across debugging sessions.
The SWE-bench family of benchmarks has expanded significantly: SWE-bench++ [31] introduces harder real-world scenarios with multi-repository dependencies, while SWE-bench-coding [32] focuses on implementation tasks rather than bug fixes. Despite these advances, even state-of-the-art models struggle with the full complexity of real-world debugging, achieving less than 25% success on these enhanced benchmarks.
@@ II-B Program Repair and Bug Localization
The debugging challenge is further highlighted by specialized benchmarks: DebugBench [35] evaluates root cause analysis, while BugHunter [36] tests multi-file bug localization. However, these benchmarks often simplify real-world debugging by providing isolated test cases rather than full repository contexts.
II-C Multi-Agent Systems for Software Tasks
The rise of multi-agent architectures has introduced new paradigms for complex software tasks. Systems like AutoGPT [37] and BabyAGI [38] demonstrate task decomposition and autonomous execution, but lack the specialized knowledge required for debugging. MetaGPT [39] simulates software company workflows with multiple specialized agents, yet still operates without persistent cross-session memory.
LangGraph [40] enables building stateful, multi-actor applications with LLMs, providing infrastructure for complex agent interactions. When combined with ReAct [41] loops, these systems can perform iterative reasoning. However, our evaluation shows that even LangGraph + ReAct configurations achieve only 31.2% debugging success compared to Chronos’s 65.3%, primarily due to: 1. Lack of persistent memory: Each debugging session starts fresh without learning from past fixes 2. Generic reasoning: No specialized understanding of debugging patterns or code dependencies 3. Inefficient exploration: Without AGR’s guided traversal, agents waste cycles on irrelevant code paths
ChatDev [42] and similar multi-agent coding systems excel at greenfield development but struggle with the complexity of debugging existing codebases. The key differentiator is that Chronos operates with continuous memory updates, specialized debugging knowledge, and efficient graph-guided retrieval, capabilities absent in generic multi-agent frameworks.
II-D How Chronos Differs
Chronos distinguishes itself from existing approaches through three fundamental architectural decisions:
1. Memory Scale: While LangChain and similar frameworks operate with session-level memory, Chronos maintains persistent debug memory (PDM) across millions of debugging sessions, learning and adapting from each interaction.
2. Debug Reasoning: Unlike generic multi-agent systems that apply general problem-solving strategies, Chronos is trained specifically on 15M+ debugging scenarios, understanding patterns like race conditions, memory leaks, and API migrations that generic models miss.
3. Orchestration: Rather than relying on user-driven or generic agent loops, Chronos implements a specialized 7-layer debugging architecture with automatic test validation, iterative refinement, and confidence-based termination.
The combination of these capabilities enables Chronos to achieve 4-5x better debugging performance than state-of-the-art alternatives, as demonstrated in our comprehensive evaluation.
II-E Performance Analysis: Code Generation vs Debugging Tasks
The landscape of code-focused LLMs has evolved dramatically in 2025. Claude Opus 4 and Sonnet 4 [4] achieve 72.5% and 72.7% respectively on SWE-bench, representing the current state-of-the-art for code generation tasks. GPT-4.1 [5] doubles GPT-4o’s performance on code diff benchmarks and reaches 54.6% on SWE-bench tasks. Gemini 2.5 Pro [30] achieves 63.8% on SWE-bench with a custom agent setup, showcasing Google’s advances in reasoning models. DeepSeek V3 [6], with 671B parameters (37B activated), demonstrates that efficient training is possible, achieving competitive performance at 1/10th the training cost of comparable models. Qwen2.5-Coder-32B [43] matches GPT-4o performance while running locally on consumer hardware.
However, these impressive code generation capabilities do not translate to debugging success. When evaluated on real-world debugging tasks requiring multi-file understanding, historical context, and iterative refinement (using COAST, MLDebugging, and MdEval benchmarks), even the best models achieve less than 15% success rates. This gap between code generation and debugging performance motivates the need for specialized debugging-focused architectures.
TABLE I: Comparison of debugging approaches across different systems and their key limitations
Approach Context Memory Debug Training Iteration Graph Key Limitation
General LLMs
Claude 4/GPT-4.1 200K-1M Session $\times$ $\times$ $\times$ No debug specialization
Gemini 2.0 Pro 2M Session $\times$ $\times$ $\times$ Attention dilution
RAG Approaches
HyDE/Self-RAG Unlimited* None $\times$ $\times$ $\times$ Static retrieval
FLARE Unlimited* None $\times$ Limited $\times$ No debug signals
Graph RAG Unlimited* Static $\times$ $\times$ Static graphs
Code Systems
SWE-agent Limited None Partial $\times$ No persistent memory
AutoCodeRover Unlimited* None Partial $\times$ Structure-only retrieval
Cursor/Windsurf Session Session $\times$ Manual $\times$ User-driven loops
Specialized
Chronos Unlimited* Persistent Auto -
*Through intelligent retrieval, not raw context window
@@ II-E Performance Analysis: Code Generation vs Debugging Tasks
III CHRONOS ARCHITECTURE: DESIGN AND IMPLEMENTATION
This section presents the technical architecture of Chronos, beginning with the fundamental insight that debugging is output-heavy rather than input-heavy, followed by the core architectural components and implementation details.
III-A Debugging as an Output-Heavy Task
@@ III-A1 Input vs Output Token Distribution
  • Total input: Often $<10K$ tokens for most real-world debugging tasks
@@ III-A2 Why Output Quality Trumps Input Size
TABLE II: Input vs output characteristics in debugging tasks.
@@
Debugging Token Flow: Input vs OutputInput (Sparse)Stack Trace: 300 tokensCode Context: 2K tokensLogs/Tests: 800 tokensHistory: 500 tokensTotal: $\sim$3.6K tokensProcessOutput (Dense)Bug Fix: 1.2K tokensExplanation: 500 tokensTests: 600 tokensDocs/PR: 400 tokensFallback: 300 tokensTotal: $\sim$3K tokensKey Insight: Output $\approx$ Input in debuggingQuality matters more than quantity
@@ III-A3 Chronos’s Output-Optimized Architecture
    Iterative Refinement Loop: Rather than single-shot generation, Chronos validates outputs through execution, using test results to refine patches, ensuring output quality over quantity.
@@
Context RetrievalMemory UsageTest IntegrationMulti-File SupportError AnalysisFix GenerationIteration SpeedCost Efficiency20%40%60%80%100%ChronosClaude 4 OpusGPT-4.1
Figure 2: Debugging capability comparison across eight key factors. Chronos (green) significantly outperforms general-purpose models in all dimensions, with particularly strong advantages in test integration (92%), iteration speed (95%), and cost efficiency (91%).
III-B High-Level Architecture: From Error Signal to Validated Fix
Kodezi Chronos is designed as an autonomous memory-driven intelligence layer for code, operating at scales that span entire enterprise repositories, team histories, and auxiliary knowledge sources. Its architecture consists of three core modules: (i) a persistent Memory Engine for continuous graph-based context construction, (ii) an advanced Retriever that constructs targeted context from code and documentation, and (iii) a transformer-based Code Reasoning Model for synthesis, debugging, and orchestration of software changes.
YesNoBug TriggeredPDMAccessAGRRetrievalChronos PlanGenerationContextAssemblyPatchDraftedTest LoopExecutedPass?Commit+ Update PDMRefine+ RetryAvg 2.2iterationsMemorypersists
Figure 3: Complete fix loop lifecycle showing integration between PDM, AGR retrieval, and iterative refinement. Dashed lines indicate feedback mechanisms that enable learning across debugging sessions.
III-C The Four Architectural Pillars
Chronos demonstrates improved debugging performance through four architectural components that distinguish it from general-purpose code models:
III-C1 Debugging-Specific Training on 15M+ Real Sessions
Unlike models trained on static code repositories, Chronos learns from:
  • 15 million debugging sessions from production environments
  • Complete fix trajectories: initial bug report $\rightarrow$ attempted fixes $\rightarrow$ test failures $\rightarrow$ successful resolution
  • Failure patterns: Common anti-patterns, regression indicators, and fix validation strategies
  • Domain-specific knowledge: Framework quirks, library-specific debugging techniques, language idioms
This specialized training enables Chronos to recognize subtle bug patterns that general models miss. For example, when encountering a React hydration mismatch, Chronos immediately knows to check server/client rendering differences rather than pursuing surface-level fixes.
III-C2 Execution Sandbox with Real-Time Feedback Loop
Chronos operates within a sophisticated execution environment that provides:
  • Isolated test execution: Every proposed fix runs in a containerized sandbox
  • Comprehensive validation: Unit tests, integration tests, linting, type checking
  • Iterative refinement: Failed fixes generate detailed error logs fed back into the next iteration
  • Regression prevention: Automatic detection of new failures introduced by fixes
This execution-driven approach explains why Chronos averages 7.8 iterations per bug while general models stop after 1-2 attempts. Each iteration refines understanding based on concrete execution results rather than probabilistic guessing.
III-C3 Persistent Repository Memory Across Sessions
Unlike stateless models that start fresh each time, Chronos maintains:
  • Bug pattern database: Historical bugs, their root causes, and successful fixes
  • Codebase evolution graph: How files, functions, and dependencies changed over time
  • Team-specific patterns: Coding conventions, common mistakes, architectural decisions
  • Dependency knowledge: Version-specific quirks, migration paths, compatibility issues
When debugging a null pointer exception, Chronos recalls similar bugs from 6 months ago, checks if recent refactoring introduced the issue, and applies team-specific null-safety patterns.
III-C4 Adaptive Graph-Guided Retrieval (AGR) for Multi-File Context
AGR enables Chronos to navigate complex codebases through:
  • Dynamic graph construction: Real-time building of dependency graphs during debugging
  • Intelligent k-hop traversal: Adaptively expanding search radius based on bug complexity, leveraging graph attention mechanisms [44]
  • Semantic + structural retrieval: Combining code semantics with architectural relationships
  • Temporal awareness: Understanding when code changed relative to bug introduction
This sophisticated retrieval explains Chronos’s 89.2% precision on the MRR benchmark, it finds the needle in the haystack by understanding the haystack’s structure.
TABLE III: Impact of each architectural pillar on debugging performance. Ablation study shows cumulative benefits.
Configuration MRR Fix Acc. SWE-bench COAST Avg. Iterations Time (min)
Base Model (no pillars) 8.3% 12.1% 9.7% 1.2 8.5
+ Debug Training 24.7% 28.3% 26.4% 2.8 16.2
+ Execution Sandbox 41.2% 43.7% 42.8% 5.1 28.4
+ Persistent Memory 55.8% 57.2% 56.3% 6.4 35.7
+ AGR (Full Chronos) 67.3% 65.3% 67.5% 7.8 42.3